# Graphite: A GPU-Accelerated Mixed-Precision Graph Optimization Framework

### v1

@@ V-B Differentiation Modes
The performance of our solver on the desktop using analytic, dynamic, and automatic differentiation modes is given in Table III for Final-13682, the largest dataset in BAL. Automatic differentiation allocates the same amount of memory as analytic differentiation, taking marginally more time at lower precisions, and significantly more at FP64. Meanwhile, the dynamic mode requires the least amount of GPU memory since Jacobians are not stored, although precision matrices are still used. While the dynamic mode exhibits a larger slowdown for FP64, this can be mitigated by switching to lower precisions. Our results show that automatic differentiation can be achieved with relatively low memory and runtime overhead for lower precisions, enabling rapid prototyping of optimizable constraints, while dynamically computed Jacobians enable more memory efficient solving at the cost of higher runtimes.
@@ V-C Case Study: Visual-Inertial Bundle Adjustment
Implementation: To use Graphite for visual-inertial bundle adjustment, it is necessary to reimplement the vertices and constraints involved in the optimization. The original data types in ORB-SLAM3 cannot be used as-is, since some contain extraneous data, involve dynamic memory allocations which must be manually transferred from the CPU to the GPU, or implement methods which are not suitable for GPU execution (e.g. a method which accesses a resource shared between different CPU threads). Therefore, we implement lightweight template data types as a replacement, which contain only the subset of data and functionality needed for visual-inertial optimization. These include data types for the combined IMU camera pose vertex, the cameras (pinhole and fisheye), and the preintegrated IMU measurements. The remaining vertex data types for the map points, velocities, gyro biases, and accelerometer biases are $3\times 1$ vectors, so we directly represent these using Eigen [30]. To use these data types with Graphite, we create five vertex descriptors.
To model the optimization constraints, we create eight descriptors for monocular and stereo reprojection errors, the inertial residual (with and without Huber loss), gyro and accelerometer bias residuals, and bias priors. The residual and analytic Jacobian expressions for each constraint are mostly a direct translation of the original ORB-SLAM3 implementation, although we found that some functionality does not behave as expected on the GPU. For example, ORB-SLAM3 normalizes a rotation matrix $R$ by first computing its SVD decomposition and then computing $R=UV^{\intercal}$, but this fails when done in a kernel, so we replace it with a Newton-Schulz iteration. For Lie groups, we use Sophus [35], as in the original implementation.
When constructing the graph, to avoid overhead from reallocating memory, we predetermine the size of the optimization, reserve memory for each descriptor, and preallocate unified memory to store the optimizable variables.
@@
Experiment: Next, large visual-inertial maps are generated by processing TUM-VI dataset sequences in ORB-SLAM3. Each map undergoes full (global) visual-inertial bundle adjustment. For the baseline, we compare our solver to the existing LDLT implementation based on g2o, using the original LM algorithm without modified termination criteria [1] for more consistent behaviour. We also include a PCG baseline, since our implementation uses PCG. For all solvers, the parameter recovery step is skipped so that all optimizations run on the same map. The optimization is set to run for a maximum of 100 iterations. Block ordering is enabled for LDLT as we found this improves the runtime performance. For g2o-PCG, we use a maximum of 50 iterations, with $1\times 10^{-6}$ absolute tolerance, while for our solver, we use a maximum of 100 iterations, with $1\times 10^{-12}$ absolute tolerance, and a rejection ratio of 10. The reason for different termination criteria is because g2o computes and solves a reduced system, instead of using the full Hessian, and we determined that this requires stricter stopping criteria for Graphite to show similar convergence. Additionally, both solvers use block Jacobi preconditioning.
@@
Results: The results for TUM-VI are given in Table IV. Note that there are approximately four vertices created for each keyframe, due to the inclusion of velocity and bias variables (e.g. a map with 3000 keyframes will have 12000 pose vertices). The LDLT baseline converges to the lowest $\chi^{2}$ error for all maps due to computing the exact parameter step at each iteration, at the cost of extremely high runtimes. Compared to the PCG baseline, our solver achieves a relative speedup of up to $59\times$ on the desktop and up to $4\times$ on the Orin Nano, despite using the full Hessian. This difference in speedup is in part due to the Orin Nano having far fewer CUDA cores and operating at lower clock speeds than its desktop counterpart (Table I). Moreover, while lower precision configurations take less time on the Jetson, double precision configurations generally achieve lower $\chi^{2}$ values and run for more iterations. Also, the Orin Nano reports higher memory usages due to the different memory architecture, so we also report the usages minus the baseline usage. Overall, Graphite achieves a significant reduction in time, enabling faster large-scale visual-inertial optimization.
@@ VI Conclusion
We have implemented a general GPU-accelerated non-linear optimization framework based on Levenberg-Marquardt for sparse estimation problems in CUDA C++, which enables the sharing of data types and functions between real-time applications and optimization tasks. Graphite achieves a relatively lower memory footprint by supporting in-place optimization, inferring matrix structures from the optimizable graph, supporting mixed-precision solving, and enabling matrix-free methods. Our results show that our solver enables large-scale optimization over complex constraints for desktop and resource-constrained devices.
Discussion: Currently, Graphite only supports Levenberg-Marquardt and PCG; direct solvers are not yet supported, so the computed parameter step is always inexact, although iterative methods scale better and use less memory. It also assumes that all constraints need observations and constraint data, but this is not always the case. Still, the overhead of storing these is mitigated by using a small data type (e.g. a char). Also, the performance could be improved further by extending vertex descriptors to support reduced camera systems, and using vectorized intrinsics and tensor cores.

### v2

@@ V-B Differentiation Modes
We report Graphite’s performance using analytic, dynamic, and automatic differentiation modes in Table II for Final-13682, the largest dataset in BAL. Automatic differentiation allocates the same amount of memory as analytic differentiation, taking approximately the same time at lower precisions, and significantly more at FP64. Meanwhile, the dynamic mode requires the least amount of GPU memory since Jacobians are not stored, although precision matrices are still used. While the dynamic mode exhibits a larger slowdown for FP64, this can be mitigated by switching to lower precisions. Our results show that automatic differentiation can be achieved with relatively low memory and runtime overhead for lower precisions, enabling rapid prototyping of optimizable constraints, while dynamically computed Jacobians enable more memory efficient solving at the cost of higher runtimes.
@@ V-C Case Study: Visual-Inertial Bundle Adjustment
Implementation: To use Graphite for visual-inertial bundle adjustment, we reimplement vertices and constraints involved in the optimization, external to the library. The original data types in ORB-SLAM3 cannot be used as-is, since some contain extraneous data, involve dynamic memory allocations which must be manually transferred from the CPU to the GPU, or implement methods which are not suitable for GPU execution (e.g. a method which accesses a resource shared between different CPU threads). Therefore, we implement lightweight template data types as a replacement, which contain only the subset of data and functionality needed for visual-inertial optimization. These include data types for the combined IMU camera pose vertex, the cameras (pinhole and fisheye), and the preintegrated IMU measurements. The remaining vertex data types for the map points, velocities, gyro biases, and accelerometer biases are $3\times 1$ vectors, so we directly represent these using Eigen [11]. To use these data types with Graphite, we define descriptors, as in 1.
To model the optimization constraints, we create eight descriptors for monocular and stereo reprojection errors, the inertial residual (with and without Huber loss), gyro and accelerometer bias residuals, and bias priors. Following the original implementation, we use the same expressions for computing the residuals and Jacobians of each constraint along with Sophus [30] for Lie groups. When constructing the graph, to avoid overhead from reallocating memory, we predetermine the size of the optimization and reserve the required memory for descriptors and optimizable variables.
@@
Experiment: Next, we evaluate the performance of visual-inertial bundle adjustment inside ORB-SLAM3 using Graphite. Among compatible datasets, we choose outdoors sequences from the TUM-VI dataset, since they generate large maps consisting of thousands of keyframes. Within ORB-SLAM3, we perform full (global) visual-inertial bundle adjustment on each of the generated maps. For the baseline, we compare Graphite to the existing LDLT-based g2o implementation, using the original LM algorithm without modified termination criteria [3] for more consistent behaviour. We also include a PCG baseline, since our implementation uses PCG. Note that the challenges discussed in Section III restrict which solvers can be used. Limiting factors include language compatibility and sufficient support for user data types involved in complex constraints (e.g. with up to six variables). For all solvers, the parameter recovery step is skipped so that all optimizations run on the same map. The optimization is set to run for up to 100 iterations. Block ordering is enabled for LDLT as we found this improves the runtime performance. For g2o PCG, we use a maximum of 50 iterations, with $1\times 10^{-6}$ absolute tolerance, while for our solver, we use a maximum of 100 iterations, with $1\times 10^{-12}$ absolute tolerance, and a rejection ratio of 10. The reason for different termination criteria is that g2o computes and solves a reduced system, while Graphite uses the full Hessian and gradient. Both solvers apply block Jacobi preconditioning.
@@
Results: The results for TUM-VI are given in Table III. The LDLT baseline converges to the lowest $\chi^{2}$ error for all maps due to computing the exact parameter step at each iteration, at the cost of extremely high runtimes. Compared to the PCG baseline, we achieve a relative speedup of up to $59\times$ on the desktop and up to $6\times$ on the Orin Nano, despite using the full Hessian. This difference in speedup is in part due to the Orin Nano having far fewer CUDA cores and operating at lower clock speeds than its desktop counterpart. Moreover, while lower precision modes take less time on the Jetson, FP64 generally achieves lower $\chi^{2}$ values and runs for more iterations. Note that the Jetson reports higher memory usages due to the different architecture, so we also include the peak usage minus the baseline usage before Graphite first runs. Overall, Graphite achieves a significant reduction in time, enabling faster large-scale visual-inertial optimization.
@@ VI Conclusion
We have implemented a general GPU-accelerated nonlinear optimization framework based on Levenberg-Marquardt for sparse estimation problems in CUDA C++, which enables the sharing of data types and functions between real-time applications and optimization tasks. Graphite achieves a relatively lower memory footprint by supporting in-place optimization, inferring matrix structures from the optimizable graph, supporting mixed-precision solving, and enabling matrix-free methods. Our results show that our framework enables large-scale optimization over complex constraints for desktop and resource-constrained devices. In the future, we would like to support more algorithms and solving techniques.